HUBBLE SPACE TELESCOPE DAILY REPORT #2153 PERIOD COVERED: 0000Z (UTC) 06/25/98 - 0000Z (UTC) 06/26/98 Daily Status Report as of 177/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Four Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.3 Completed NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxy NGC-891 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.4 Completed Two Sets of NIC/1/2/3 7448 (Star Formation in a Protogalaxy Candidate at z=2.72) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the star formation region within the protogalaxy candidate MS1512+36 which is located at a redshift of Z = 2.72 . Recently a new population of galaxies at high redshift has been identified by ultraviolet absorption lines which most likely arise from very young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.5 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 8049 (Long WF/PC-2 Decontaminations during 2nd NIC/3 Campaign) The WF/PC-2 was used to make long decontamination procedures as part of the second NIC/3 campaign. Each decon is followed by five 30-minute darks. During this time, no external science WF/PC-2 observations were executed. The decontaminations were executed as scheduled, and no anomalies were noted. 1.7 Completed NIC/3 7841 (HST/NICMOS Observations of the Nearest Embedded Cluster: the L1688 {Rho Oph} Protostellar Region) The Near Infrared Camera (Camera Three) was used to make observations protostellar region of the nearest young cluster L1688 (Positions 8, 9 and 10) ( a.k.a., Rho Oph). This cluster is unique in its proximity to Earth, only 125 pc away; and in its high proportion of optically invisible protostars. The observations were executed as scheduled, and no anomalies were noted. .0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. NICMOS flight software (FSW) 4.0B was successfully installed at 176/1818Z. This enhancement to FSW 4.0A corrects the function which determines the position of the coronagraphic hole. It also adds a new status buffer message (NICMOS 740) which will indicate if the calculated centroid checkbox was outside the field of regard for the hole. A real-time FHST map was commanded at 177/0021Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The NICMOS Camera Three Campaign will be ending this Sunday evening. Upon its conclusion, the secondary mirror will be returned to its nominal focus position, a move of approximately -15.2 microns. The earliest time for the mirror move is 179/1725Z. Continuation of normal science observations and calibrations. /DMH